A solid sphere of radius 5 cm floats in water.If a maximun load of 0.1 kg can be put on it without wetting the load, find the specific gravity of the material of the sphere???

Grade:11

1 Answers

Swapnil Saxena
12 years ago

Considering the equilibrium state after the load is put in the water.

We can say that

Mg (M is the mass of the sphere) + mg (m is the mass of the load)= F(buoyancy)

Mg (M is the mass of the sphere) + mg (m is the mass of the load)= Vpg (V  is the part of the system under water=Volume of Sphere)

M+m= Vp

M/V(density of sphere) = (p-m/V)

Specific gravity = M/pV= 1-m/pV

m= 0.1

V=4/3pi(5*10-2)3

p=103(density of water)

So value of specific gravity can be found